Circular No. 6020

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

PERIODIC COMET SHOEMAKER-LEVY 9 (1993e) Z. Sekanina, Jet Propulsion Laboratory, writes: “Preliminary results

of analysis of CCD photometry obtained with the Hubble Space Telescope’s Planetary Camera in late January, late March, and early July suggest that the deconvolved signal for the largest fragment in each nuclear condensation does not appear to depend strongly on the law employed to subtract the contribution from the dust cloud within ∼ 0′′.35 (equivalent to ∼ 1350 km in late January, 1100 km in late March, and 1200 km in early July) around the brightest pixel. Two different laws employed for the brightness distribution of the extended source, ∼ exp[−(ρ^{2}/2σ^{2})^{ν}] and ∼ [1 + (ρ/σ)^{ν}]^{−1} (where ρ is the angular distance from the source’s center and σ and ν are parametric constants), imply the presence of a major point source [defined as a body that is characterized by a single point-spread function (PSF)] in the center of most of the condensations and offer consistent values for the dimensions of the largest fragments, which for an assumed geometrical albedo of 4 percent in the R filter and a phase coefficient of 0.035 mag/deg yield the effective diameters listed below. In addition, most of the brightness profiles imply the presence of isolated clumps of slightly elevated signals, which fit the PSF’s profile and may represent companions that recently split off from the parent fragment. Most of them may break up further before they can be detected from the ground. Their positions are found to be independent of the brightness law assumed for the extended source, but their numbers and configurations vary with time. Their signals suggest effective diameters generally between 1 and 2 km (∼ 1 km is the smallest size that the HST can individually detect at the comet’s distance). Following are listed the fragment designation, its largest effective diameter (km) in Jan., Mar., and July, and the number of possible companions detected during the same times: F = 16, 2.3, –, –, 1, –, –; G = 15, 4.2, 3.6, –, 2–4, 6–8, –; H = 14, 3.2, –, –, 2, –, –; P_{1} = 8a, 1.3, <1, –, 2, –, –; P_{2} = 8b, 2.4, –, –, 3, –, –; Q_{1} = 7a, 3.9, 3.0, 3.7, 2–5, 2–3, 3–5; Q_{2} = 7b, 3.2, 1.6, 2.4, 2, 2–4, 1–2; R = 6, 2.7, 2.2, –, 0, 2, –; S = 5, 3.6, –, –, 6, –, –; T = 4, 1.4, –, –, 1, –, –; U = 3, 1.3, –, –, 0, –, –. The peak of the dust cloud is found to be displaced with respect to the primary fragment in the general direction away from the sun. Although the evidence points to an apparently continuing disintegration of the large fragments in numerous discrete events, objects a few km across still seem to have been present in early July, and the temporal variations in the effective diameters are likely to be primarily a rotational effect of strongly irregular shape.”

1994 July 15 Daniel W. E. Green